# One Fits All: Learning Fair Graph Neural Networks for Various Sensitive Attributes

### v1

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3672029isbn: 979-8-4007-0490-1/24/08isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Knowledge representation and reasoning
@@ 1. Introduction
Graph neural networks (GNNs) have achieved tremendous success in processing graph-structured data, such as citation networks (Tang et al., 2008; Shchur et al., 2018) and social networks (Leskovec and Mcauley, 2012; Takac and Zabovsky, 2012). Consequently, this advancement has led to their application across diverse domains, including fraud detection (Dou et al., 2020), recommender systems (Wu et al., 2022a), and drug discovery (Bongini et al., 2021). However, recent studies (Dai and Wang, 2021; Spinelli et al., 2021) have unveiled a concerning trend that GNNs make discriminatory decisions toward the specific protected groups defined by sensitive attributes, e.g., race, and age. This phenomenon, termed the group fairness problem of GNNs, hinders the application of GNNs in high-stake scenarios.
@@ 2.1. Fairness in Graph Neural Networks
The fairness of GNNs includes group fairness (Dai and Wang, 2021; Agarwal et al., 2021) and individual fairness (Kang et al., 2020; Dong et al., 2021). Our study focuses on the group fairness aspect, emphasizing equitable model decisions for each protected group partitioned by the sensitive attribute. Recent studies improving group fairness in GNNs typically segregate into pre-process (Dong et al., 2022; Ling et al., 2022) and in-process (Wang et al., 2022; Bose and Hamilton, 2019) approaches. Pre-process approaches aim to mitigate biases in training data before training downstream tasks. To mitigate biases, techniques like adversarial learning (Ling et al., 2022), and distribution alignment (Dong et al., 2022; Current et al., 2022; Yang et al., 2022) serve as optimization objectives for debiasing training data. Additionally, some heuristic approaches modify the training graph (Spinelli et al., 2021; Li et al., 2022a) or reweight edge (Khajehnejad et al., 2022; Li et al., 2021) by either enhancing connections between diverse groups or reducing connections within the same groups. In-process approaches aim to train fair GNNs through the fairness-aware framework. Similar to pre-process approaches, in-process approaches also incorporate adversarial learning (Bose and Hamilton, 2019; Dai and Wang, 2021) and distribution alignment (Guo et al., 2022; Fan et al., 2021) to learn GNNs. Despite significant progress, these approaches are tailored to the specific sensitive attribute, lacking considerations for various sensitive attributes.

### v2

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3672029isbn: 979-8-4007-0490-1/24/08ccs: Computing methodologies Knowledge representation and reasoning
@@ 1. Introduction
Graph neural networks (GNNs) have achieved tremendous success in processing graph-structured data (Kipf and Welling, 2016; Li et al., 2023, 2024), such as citation networks (Shchur et al., 2018) and social networks (Leskovec and Mcauley, 2012; Takac and Zabovsky, 2012). Consequently, this advancement has led to their application across diverse domains, including fraud detection (Dou et al., 2020) and recommender systems (Wu et al., 2022a). However, recent studies (Dai and Wang, 2021; Spinelli et al., 2021) have unveiled a concerning trend that GNNs make discriminatory decisions toward the specific protected groups defined by sensitive attributes, e.g., race, and age. This phenomenon, termed the group fairness problem of GNNs, hinders the application of GNNs in high-stake scenarios.
@@ 2.1. Fairness in Graph Neural Networks
The fairness of GNNs includes group fairness (Dai and Wang, 2021; Agarwal et al., 2021; Zhu et al., 2024c) and individual fairness (Kang et al., 2020; Dong et al., 2021). Our study focuses on the group fairness aspect, emphasizing equitable model decisions for each protected group partitioned by the sensitive attribute. Recent studies improving group fairness in GNNs typically segregate into pre-process (Dong et al., 2022; Ling et al., 2022; Zhu et al., 2024b) and in-process (Wang et al., 2022; Bose and Hamilton, 2019; Zhu et al., 2024a) approaches. Pre-process approaches aim to mitigate biases in training data before training downstream tasks. To mitigate biases, techniques like adversarial learning (Ling et al., 2022), and distribution alignment (Dong et al., 2022; Current et al., 2022; Yang et al., 2022) serve as optimization objectives for debiasing training data. Additionally, some heuristic approaches modify the training graph (Spinelli et al., 2021; Li et al., 2022a) or reweight edge (Khajehnejad et al., 2022; Li et al., 2021) by either enhancing connections between diverse groups or reducing connections within the same groups. In-process approaches aim to train fair GNNs through the fairness-aware framework. Similar to pre-process approaches, in-process approaches also incorporate adversarial learning (Bose and Hamilton, 2019; Dai and Wang, 2021) and distribution alignment (Guo et al., 2022; Fan et al., 2021) to learn GNNs. Despite significant progress, these approaches are tailored to the specific sensitive attribute, lacking considerations for various sensitive attributes.